OpenStudy (tiffany_rhodes):

If:

OpenStudy (tiffany_rhodes):

\[\int\limits_{-3}^{0} f(x)dx = 2 \] and \[\int\limits_{2}^{3} g(x)dx = -4\] what is the value of \[\int\limits_{?}^{?}\int\limits_{?}^{?} f(x)g(y) dA\] where D is the square:

OpenStudy (tiffany_rhodes):

\[-3\le x \le 0\]

OpenStudy (tiffany_rhodes):

\[2 \le y \le 3\]

jimthompson5910 (jim_thompson5910):

it depends on the order of integration. If you integrate x first, then the limits are from -3 to 0 and they apply to the inner most integral \[\large \int\limits_{?}^{?}\int\limits_{-3}^{0} f(x)g(y) dA\]

jimthompson5910 (jim_thompson5910):

the limits of integration for y are applied to the outer integral \[\large \int\limits_{2}^{3}\int\limits_{-3}^{0} f(x)g(y) dA\]

jimthompson5910 (jim_thompson5910):

and you change the dA to dxdy the dx comes first to say you integrate wrt x first, then y comes next \[\large \int\limits_{2}^{3}\int\limits_{-3}^{0} f(x)g(y)dxdy\]

jimthompson5910 (jim_thompson5910):

hopefully you see how to finish up

OpenStudy (tiffany_rhodes):

I'm not seeing what function(s) to integrate . . .

jimthompson5910 (jim_thompson5910):

It might help to pull out g(y). This entire function is a constant because g(y) has no x's present in it. It's a function of y. That allows you to do this \[\large \int\limits_{2}^{3}\int\limits_{-3}^{0} f(x)g(y)dxdy\] \[\large \int\limits_{2}^{3}g(y)\left(\int\limits_{-3}^{0} f(x)dx\right)dy\]

jimthompson5910 (jim_thompson5910):

you can't pull g(y) out entirely because g(y) is no longer a constant for the outer integral (wrt y)

OpenStudy (tiffany_rhodes):

Okay so when you take the antiderivative of g(y) would it just be y? or would it be (y^2)/2 . . . I'm unsure what function I'm actually integrating.

jimthompson5910 (jim_thompson5910):

don't worry about the outer integral right now

jimthompson5910 (jim_thompson5910):

focus on the inner integral \[\large \int\limits_{-3}^{0} f(x)dx\]

jimthompson5910 (jim_thompson5910):

normally you'd integrate using rules you learn in calc 2, but you're not given the definition of f(x). However, they do provide the value of the definite integral above.

OpenStudy (tiffany_rhodes):

okay, so the value is 2

jimthompson5910 (jim_thompson5910):

correct, allowing you to say \[\large \int\limits_{2}^{3}g(y)\left(\int\limits_{-3}^{0} f(x)dx\right)dy = \int\limits_{2}^{3}g(y)\left(2\right)dy\]

jimthompson5910 (jim_thompson5910):

then the 2 is a constant which you can pull out to get \[\large 2\int\limits_{2}^{3}g(y)dy\]

OpenStudy (tiffany_rhodes):

alright

OpenStudy (tiffany_rhodes):

Ooh, okay. I think I see where this problem is going. Basically you can multiply the two values of the integrals together to get the answer?

jimthompson5910 (jim_thompson5910):

yeah pretty much, so \[\large \int\limits_{2}^{3}\int\limits_{-3}^{0} f(x)g(y)dxdy = -8\] which means \[\large \int\limits\int f(x)g(y)dA = -8\] where A is the square created by \(-3 \le x \le 0\) and \(2 \le y \le 3\)

jimthompson5910 (jim_thompson5910):

And you can reverse the order of integration if you wanted. You could start off integrating with respect to y first, then move onto x. You'd get the same answer either way.

OpenStudy (tiffany_rhodes):

okay, thanks for the help and sorry about the confusion

jimthompson5910 (jim_thompson5910):

but you would have dydx if you did that

jimthompson5910 (jim_thompson5910):

that's ok